OpenStudy (anonymous):

Use the elimination method to solve the following system of equations. 2x – y + z = –3 2x + 2y + 3z = 2 3x – 3y – z = –4

OpenStudy (anonymous):

Solve 3 x-3 y-z = -4 (equation 1) 2 x+2 y+3 z = 2 (equation 2) 2 x-y+z = -3 (equation 3) Subtract 2/3 x (equation 1) from equation 2: 3 x-3 y-z = -4 (equation 1) 0 x+4 y+(11 z)/3 = 14/3 (equation 2) 2 x-y+z = -3 (equation 3) Multiply equation 2 by 3: 3 x-3 y-z = -4 (equation 1) 0 x+12 y+11 z = 14 (equation 2) 2 x-y+z = -3 (equation 3) Subtract 2/3 × (equation 1) from equation 3: 3 x-3 y-z = -4 | (equation 1) 0 x+12 y+11 z = 14 | (equation 2) 0 x+y+(5 z)/3 = -1/3 | (equation 3) Multiply equation 3 by 3: 3 x-3 y-z = -4 | (equation 1) 0 x+12 y+11 z = 14 | (equation 2) 0 x+3 y+5 z = -1 | (equation 3) Subtract 1/4 × (equation 2) from equation 3: 3 x-3 y-z = -4 (equation 1) 0 x+12 y+11 z = 14 (equation 2) 0 x+0 y+(9 z)/4 = -9/2 (equation 3) Multiply equation 3 by 4/9: 3 x-3 y-z = -4 (equation 1) 0 x+12 y+11 z = 14 (equation 2) 0 x+0 y+z = -2 (equation 3) Subtract 11 × (equation 3) from equation 2: 3 x-3 y-z = -4 (equation 1) 0 x+12 y+0 z = 36 (equation 2) 0 x+0 y+z = -2 (equation 3) Divide equation 2 by 12: 3 x-3 y-z = -4 (equation 1) 0 x+y+0 z = 3 (equation 2) 0 x+0 y+z = -2 (equation 3) Add 3 × (equation 2) to equation 1: 3 x+0 y-z = 5 (equation 1) 0 x+y+0 z = 3 (equation 2) 0 x+0 y+z = -2 (equation 3) Add equation 3 to equation 1: 3 x+0 y+0 z = 3 (equation 1) 0 x+y+0 z = 3 (equation 2) 0 x+0 y+z = -2 (equation 3) Divide equation 1 by 3: {x+0 y+0 z = 1 (equation 1) 0 x+y+0 z = 3 (equation 2) 0 x+0 y+z = -2 (equation 3) Answer: x = 1 , y = 3 , z = -2

OpenStudy (anonymous):

thank you so muchhhh

OpenStudy (anonymous):

no problem :)